OpenStudy (anonymous):

Did I expand it wrong? y= (1+cos^2 (x))^6 y' = 6 ( 1+cos^2 (x) )^5 (2) (cosx) (-sinx)

OpenStudy (shayaan_mustafa):

Hi cmkc109 :) No I don't think so. It seems right buddy..

OpenStudy (anonymous):

But then the answer given is -6 sin 2x (1+cos^2 (x))^5

OpenStudy (shayaan_mustafa):

sin(2x)?? double angle?

OpenStudy (shayaan_mustafa):

OK let me try..

OpenStudy (anonymous):

it doesnt have the bracket..but i assume it is the same -6 sin 2x (1+cos^2 x)^5

OpenStudy (anonymous):

since 2sinxcosx = sin2x

OpenStudy (anonymous):

That's all you do to simply further!

OpenStudy (anonymous):

ohh..that rule, thanks both of you :)

OpenStudy (shayaan_mustafa):

\[\large y=(1+\cos^2x)^6\]\[\large y'=6(1+\cos^2x)^5\frac{d}{dx}(1+\cos^2x)\]\[\large y'=6(1+cos^2x)^5(2cosx)(-sinx)\]\[\large y'=-6(1+cos^2x)^5(2sinxcosx)\]\[\large y'=-6(1+cos^2x)^5(sin2x)\]

OpenStudy (anonymous):

thanks both!